# MeteoRA: Multiple-tasks Embedded LoRA for Large Language Models

### v2

@@ A.4 Experimental results of 28 tasks
| Task Name | PEFT (reference) | MeteoRA (top-1) | MeteoRA (top-2) | LoRA-F | LoRA-B |
| abstract_narrative_understanding | 0.803 | 0.787 | 0.787 | 0.793 | 0.790 |
| contextual_parametric_knowledge_conflicts | 0.999 | 0.976 | 0.999 | 0.999 | 0.999 |
| cs_algorithms | 0.841 | 0.826 | 0.845 | 0.841 | 0.841 |
| disfl_qa | 0.703 | 0.703 | 0.706 | 0.680 | 0.605 |
| elementary_math_qa | 0.780 | 0.773 | 0.776 | 0.777 | 0.606 |
| epistemic_reasoning | 1.000 | 1.000 | 1.000 | 0.996 | 1.000 |
| formal_fallacies_syllogisms_negation | 0.989 | 0.987 | 0.987 | 0.970 | 0.628 |
| goal_step_wikihow | 0.935 | 0.928 | 0.932 | 0.921 | 0.923 |
| gsm8k | 0.591 | 0.559 | 0.555 | 0.566 | 0.548 |
| language_identification | 0.782 | 0.775 | 0.779 | 0.749 | 0.649 |
| logical_deduction | 0.760 | 0.753 | 0.757 | 0.707 | 0.707 |
| object_counting | 0.880 | 0.850 | 0.875 | 0.555 | 0.865 |
| paragraph_segmentation | 0.296 | 0.252 | 0.295 | 0.261 | 0.244 |
| play_dialog_same_or_different | 0.649 | 0.649 | 0.649 | 0.632 | 0.649 |
| question_selection | 0.936 | 0.940 | 0.927 | 0.911 | 0.930 |
| reasoning_about_colored_objects | 0.958 | 0.943 | 0.960 | 0.945 | 0.965 |
| strategyqa | 0.716 | 0.670 | 0.659 | 0.707 | 0.718 |
| tracking_shuffled_objects | 0.995 | 0.996 | 0.993 | 0.588 | 0.664 |
| unit_conversion | 0.822 | 0.819 | 0.820 | 0.814 | 0.780 |
| vitaminc_fact_verification | 0.908 | 0.907 | 0.907 | 0.903 | 0.839 |
| winowhy | 0.816 | 0.827 | 0.818 | 0.797 | 0.802 |
| Average | 0.817 | 0.806 | 0.811 | 0.767 | 0.750 |
@@
| cnn_dailymail | PEFT (reference) | 7.50 | 0.228 | 0.067 | 0.214 |
| MeteoRA (top-1) | 7.50 | 0.229 | 0.069 | 0.216 | |
| MeteoRA (top-2) | 5.57 | 0.230 | 0.070 | 0.217 | |
@@
| linguistics_puzzles | PEFT (reference) | 46.17 | 0.716 | 0.479 | 0.659 |
| MeteoRA (top-1) | 68.34 | 0.717 | 0.478 | 0.661 | |
| MeteoRA (top-2) | 46.17 | 0.713 | 0.476 | 0.655 | |
@@
| news_commentary_de | PEFT (reference) | 78.25 | - | - | - |
| MeteoRA (top-1) | 86.48 | - | - | - | |
| MeteoRA (top-2) | 86.48 | - | - | - | |
@@
| news_commentary_es | PEFT (reference) | 70.05 | - | - | - |
| MeteoRA (top-1) | 81.54 | - | - | - | |
| MeteoRA (top-2) | 70.05 | - | - | - | |
@@
| news_commentary_it | PEFT (reference) | 39.04 | - | - | - |
| MeteoRA (top-1) | 39.04 | - | - | - | |
| MeteoRA (top-2) | 39.04 | - | - | - | |
@@
| strategyqa | PEFT (reference) | 15.72 | 0.237 | 0.064 | 0.222 |
| MeteoRA (top-1) | 8.74 | 0.235 | 0.065 | 0.221 | |
| MeteoRA (top-2) | 10.03 | 0.240 | 0.068 | 0.226 | |
@@
| topical_chat | PEFT (reference) | 12.50 | 0.157 | 0.027 | 0.146 |
| MeteoRA (top-1) | 13.44 | 0.151 | 0.025 | 0.141 | |
| MeteoRA (top-2) | 12.35 | 0.149 | 0.025 | 0.140 | |
@@
| LoRA-B | 9.78 | 0.143 | 0.021 | 0.134 |
@@
| cnn_dailymail | PEFT (reference) | 11.93 | 0.231 | 0.069 | 0.218 |
| MeteoRA (top-1) | 11.93 | 0.232 | 0.070 | 0.219 | |
| MeteoRA (top-2) | 11.93 | 0.233 | 0.070 | 0.218 | |
@@
| linguistics_puzzles | PEFT (reference) | 41.72 | 0.697 | 0.460 | 0.638 |
| MeteoRA (top-1) | 41.72 | 0.696 | 0.448 | 0.639 | |
| MeteoRA (top-2) | 41.72 | 0.695 | 0.451 | 0.636 | |
@@
| news_commentary_de | PEFT (reference) | 97.65 | - | - | - |
| MeteoRA (top-1) | 86.48 | - | - | - | |
| MeteoRA (top-2) | 86.48 | - | - | - | |
@@
| news_commentary_es | PEFT (reference) | 81.54 | - | - | - |
| MeteoRA (top-1) | 81.54 | - | - | - | |
| MeteoRA (top-2) | 81.54 | - | - | - | |
@@
| news_commentary_it | PEFT (reference) | 54.90 | - | - | - |
| MeteoRA (top-1) | 51.83 | - | - | - | |
| MeteoRA (top-2) | 54.90 | - | - | - | |
@@
| strategyqa | PEFT (reference) | 10.58 | 0.249 | 0.077 | 0.236 |
| MeteoRA (top-1) | 10.58 | 0.250 | 0.077 | 0.239 | |
| MeteoRA (top-2) | 10.58 | 0.252 | 0.076 | 0.239 | |
@@
| topical_chat | PEFT (reference) | 39.50 | 0.151 | 0.025 | 0.141 |
| MeteoRA (top-1) | 45.64 | 0.152 | 0.024 | 0.141 | |
| MeteoRA (top-2) | 45.64 | 0.152 | 0.026 | 0.141 | |
@@
| LoRA-B | 19.22 | 0.139 | 0.019 | 0.131 |
@@ A.5 Composite-n evaluation results details
| logical_deduction | 0.500 | 0.430 | - | - | - | - | - | - | - | - |
| question_selection | 0.545 | 0.630 | - | - | - | - | - | - | - | - |
| strategyqa | 0.445 | 0.250 | 15.31 | 10.55 | 0.195 | 0.135 | 0.052 | 0.027 | 0.182 | 0.128 |
@@
| logical_deduction | 0.500 | 0.500 | - | - | - | - | - | - | - | - |
| question_selection | 0.620 | 0.770 | - | - | - | - | - | - | - | - |
| abstract_narrative_understanding | 0.350 | 0.460 | - | - | - | - | - | - | - | - |
| goal_step_wikihow | 0.650 | 0.410 | - | - | - | - | - | - | - | - |
| strategyqa | 0.495 | 0.275 | 9.86 | 9.41 | 0.221 | 0.219 | 0.069 | 0.063 | 0.207 | 0.208 |
@@
| logical_deduction | 0.500 | 0.453 | - | - | - | - | - | - | - | - |
| question_selection | 0.703 | 0.688 | - | - | - | - | - | - | - | - |
| abstract_narrative_understanding | 0.625 | 0.672 | - | - | - | - | - | - | - | - |
| goal_step_wikihow | 0.773 | 0.727 | - | - | - | - | - | - | - | - |
@@
| strategyqa | 0.461 | 0.211 | 3.23 | 0.00 | 0.225 | 0.106 | 0.051 | 0.025 | 0.210 | 0.099 |
| disfl_qa | 0.266 | 0.117 | - | - | - | - | - | - | - | - |
| news_commentary_de | - | - | 14.78 | 14.54 | - | - | - | - | - | - |

### v3

@@ A.4 Experimental results of 28 tasks
Task Name | PEFT (reference) | LORA-F | LORA-B | Avg LoRA | TIES | DARE | Arrow | LoraHub | MeteoRA (T1-1k) | MeteoRA (T2-1k) | MeteoRA(T2-100) | MeteoRA(T2-5) |
AbsNarr | 0.803 | 0.793 | 0.790 | 0.413 | 0.425 | 0.335 | 0.772 | 0.075 | 0.787 | 0.787 | 0.775 | 0.768 |
ConParaKC | 0.999 | 0.999 | 0.999 | 0.514 | 0.594 | 0.492 | 0.997 | 0.219 | 0.999 | 0.999 | 0.976 | 0.992 |
CSAlg | 0.841 | 0.841 | 0.841 | 0.705 | 0.686 | 0.663 | 0.780 | 0.602 | 0.845 | 0.826 | 0.826 | 0.830 |
DisflQA | 0.703 | 0.680 | 0.605 | 0.374 | 0.396 | 0.377 | 0.504 | 0.197 | 0.706 | 0.703 | 0.686 | 0.628 |
ElemMath | 0.780 | 0.777 | 0.606 | 0.273 | 0.308 | 0.245 | 0.645 | 0.106 | 0.776 | 0.773 | 0.751 | 0.725 |
EpiReason | 1.000 | 0.996 | 1.000 | 0.430 | 0.450 | 0.425 | 0.600 | 0.170 | 1.000 | 1.000 | 1.000 | 1.000 |
FormFall | 0.989 | 0.970 | 0.628 | 0.528 | 0.519 | 0.520 | 0.836 | 0.190 | 0.987 | 0.987 | 0.981 | 0.977 |
GSWiki | 0.935 | 0.921 | 0.923 | 0.627 | 0.608 | 0.574 | 0.835 | 0.307 | 0.932 | 0.928 | 0.904 | 0.896 |
GSM8K | 0.591 | 0.566 | 0.548 | 0.080 | 0.086 | 0.108 | 0.172 | 0.050 | 0.555 | 0.559 | 0.511 | 0.491 |
LangID | 0.782 | 0.749 | 0.649 | 0.404 | 0.412 | 0.383 | 0.625 | 0.192 | 0.779 | 0.775 | 0.759 | 0.744 |
LogDeduc | 0.760 | 0.707 | 0.707 | 0.403 | 0.423 | 0.383 | 0.627 | 0.367 | 0.757 | 0.753 | 0.747 | 0.770 |
ObjCount | 0.880 | 0.555 | 0.865 | 0.060 | 0.080 | 0.130 | 0.005 | 0.230 | 0.875 | 0.850 | 0.785 | 0.750 |
ParaSeg | 0.296 | 0.261 | 0.244 | 0.044 | 0.050 | 0.045 | 0.187 | 0.000 | 0.295 | 0.252 | 0.235 | 0.234 |
PlayDiag | 0.649 | 0.632 | 0.649 | 0.647 | 0.650 | 0.644 | 0.656 | 0.092 | 0.649 | 0.649 | 0.580 | 0.581 |
QuestSel | 0.936 | 0.911 | 0.930 | 0.544 | 0.506 | 0.472 | 0.845 | 0.247 | 0.927 | 0.940 | 0.892 | 0.892 |
ColorReason | 0.958 | 0.945 | 0.965 | 0.565 | 0.595 | 0.530 | 0.793 | 0.238 | 0.960 | 0.983 | 0.915 | 0.905 |
StratQA | 0.716 | 0.707 | 0.718 | 0.600 | 0.611 | 0.538 | 0.681 | 0.503 | 0.659 | 0.670 | 0.648 | 0.611 |
TrackObj | 0.995 | 0.588 | 0.664 | 0.147 | 0.195 | 0.136 | 0.804 | 0.171 | 0.993 | 0.996 | 0.985 | 0.985 |
UnitConv | 0.822 | 0.814 | 0.780 | 0.485 | 0.491 | 0.410 | 0.647 | 0.463 | 0.820 | 0.819 | 0.802 | 0.786 |
VitaFact | 0.908 | 0.903 | 0.839 | 0.607 | 0.655 | 0.541 | 0.822 | 0.311 | 0.907 | 0.907 | 0.902 | 0.890 |
WinoWhy | 0.816 | 0.797 | 0.802 | 0.524 | 0.516 | 0.526 | 0.750 | 0.203 | 0.818 | 0.827 | 0.788 | 0.788 |
Average | 0.817 | 0.767 | 0.750 | 0.427 | 0.441 | 0.404 | 0.647 | 0.235 | 0.811 | 0.806 | 0.783 | 0.773 |
@@
Avg LoRA | 41.88 | 0.195 | 0.084 | 0.164 | |
TIES | 80.34 | 0.209 | 0.092 | 0.175 | |
DARE | 78.25 | 0.228 | 0.101 | 0.193 | |
Arrow | 24.62 | 0.271 | 0.128 | 0.230 | |
LoraHub | 0.00 | 0.240 | 0.117 | 0.206 | |
MeteoRA (T1-1k) | 28.83 | 0.350 | 0.166 | 0.329 | |
MeteoRA (T2-1k) | 24.12 | 0.349 | 0.162 | 0.327 | |
MeteoRA (T2-100) | 39.09 | 0.332 | 0.160 | 0.281 | |
MeteoRA (T2-5) | 12.49 | 0.306 | 0.140 | 0.256 | |
CNNDM | PEFT (reference) | 7.50 | 0.228 | 0.067 | 0.214 |
@@
Avg LoRA | 13.08 | 0.144 | 0.032 | 0.104 | |
TIES | 13.08 | 0.147 | 0.032 | 0.104 | |
DARE | 13.08 | 0.126 | 0.031 | 0.081 | |
Arrow | 17.42 | 0.173 | 0.043 | 0.122 | |
LoraHub | 4.77 | 0.141 | 0.030 | 0.104 | |
MeteoRA (T1-1k) | 7.50 | 0.229 | 0.069 | 0.216 | |
MeteoRA (T2-1k) | 5.57 | 0.230 | 0.070 | 0.217 | |
MeteoRA (T2-100) | 7.32 | 0.251 | 0.070 | 0.196 | |
MeteoRA (T2-5) | 7.77 | 0.254 | 0.073 | 0.199 | |
LingPuzz | PEFT (reference) | 46.17 | 0.716 | 0.479 | 0.659 |
@@
Avg Lora | 36.72 | 0.531 | 0.233 | 0.441 | |
TIES | 49.14 | 0.405 | 0.117 | 0.308 | |
DARE | 68.87 | 0.379 | 0.102 | 0.285 | |
Arrow | 56.23 | 0.643 | 0.365 | 0.562 | |
LoraHub | 0.00 | 0.172 | 0.057 | 0.131 | |
MeteoRA (T1-1k) | 68.34 | 0.717 | 0.478 | 0.661 | |
MeteoRA (T2-1k) | 46.17 | 0.713 | 0.476 | 0.655 | |
MeteoRA (T2-100) | 46.17 | 0.718 | 0.480 | 0.646 | |
MeteoRA (T2-5) | 57.47 | 0.716 | 0.474 | 0.646 | |
NewsDE | PEFT (reference) | 78.25 | - | - | - |
@@
Avg Lora | 3.38 | - | - | - | |
TIES | 86.48 | - | - | - | |
DARE | 86.48 | - | - | - | |
Arrow | 86.48 | - | - | - | |
LoraHub | 50.09 | - | - | - | |
MeteoRA (T1-1k) | 86.48 | - | - | - | |
MeteoRA (T2-1k) | 86.48 | - | - | - | |
MeteoRA (T2-100) | 86.48 | - | - | - | |
MeteoRA (T2-5) | 86.48 | - | - | - | |
NewsES | PEFT (reference) | 70.05 | - | - | - |
@@
Avg Lora | 2.86 | - | - | - | |
TIES | 70.05 | - | - | - | |
DARE | 46.27 | - | - | - | |
Arrow | 81.54 | - | - | - | |
LoraHub | 0.64 | - | - | - | |
MeteoRA (T1-1k) | 81.54 | - | - | - | |
MeteoRA (T2-1k) | 70.05 | - | - | - | |
MeteoRA (T2-100) | 70.05 | - | - | - | |
MeteoRA (T2-5) | 70.05 | - | - | - | |
NewsIT | PEFT (reference) | 39.04 | - | - | - |
@@
Avg Lora | 40.20 | - | - | - | |
TIES | 40.08 | - | - | - | |
DARE | 40.20 | - | - | - | |
Arrow | 36.92 | - | - | - | |
LoraHub | 40.08 | - | - | - | |
MeteoRA (T1-1k) | 39.04 | - | - | - | |
MeteoRA (T2-1k) | 39.04 | - | - | - | |
MeteoRA (T2-100) | 39.04 | - | - | - | |
MeteoRA (T2-1k) | 39.04 | - | - | - | |
StratQA | PEFT (reference) | 15.72 | 0.237 | 0.064 | 0.222 |
@@
Avg Lora | 14.54 | 0.185 | 0.050 | 0.149 | |
TIES | 16.62 | 0.112 | 0.024 | 0.088 | |
DARE | 16.62 | 0.123 | 0.026 | 0.095 | |
Arrow | 13.83 | 0.218 | 0.066 | 0.175 | |
LoraHub | 11.50 | 0.171 | 0.038 | 0.128 | |
MeteoRA (T1-1k) | 8.74 | 0.235 | 0.065 | 0.221 | |
MeteoRA (T2-1k) | 10.03 | 0.240 | 0.068 | 0.226 | |
MeteoRA (T2-100) | 13.95 | 0.222 | 0.063 | 0.172 | |
MeteoRA (T2-5) | 20.69 | 0.228 | 0.067 | 0.174 | |
TopChat | PEFT (reference) | 12.50 | 0.157 | 0.027 | 0.146 |
@@
LoRA-B | 9.78 | 0.143 | 0.021 | 0.134 | |
Avg Lora | 1.21 | 0.099 | 0.010 | 0.060 | |
TIES | 22.45 | 0.101 | 0.011 | 0.078 | |
DARE | 22.48 | 0.103 | 0.011 | 0.065 | |
Arrow | 11.16 | 0.099 | 0.013 | 0.080 | |
LoraHub | 0.35 | 0.064 | 0.005 | 0.051 | |
MeteoRA (T1-1k) | 13.44 | 0.151 | 0.025 | 0.141 | |
MeteoRA (T2-1k) | 12.35 | 0.149 | 0.025 | 0.140 | |
MeteoRA (T2-100) | 13.44 | 0.132 | 0.023 | 0.108 | |
MeteoRA (T2-5) | 12.93 | 0.135 | 0.024 | 0.110 |
@@
Avg LoRA | 73.49 | 0.206 | 0.089 | 0.172 | |
TIES | 73.49 | 0.214 | 0.092 | 0.181 | |
DARE | 73.49 | 0.230 | 0.099 | 0.192 | |
Arrow | 12.26 | 0.222 | 0.093 | 0.186 | |
LoraHub | 0.00 | 0.176 | 0.068 | 0.151 | |
MeteoRA (T1-1k) | 32.34 | 0.358 | 0.170 | 0.335 | |
MeteoRA (T2-1k) | 30.08 | 0.354 | 0.170 | 0.332 | |
MeteoRA (T2-100) | 31.19 | 0.317 | 0.147 | 0.266 | |
MeteoRA (T2-5) | 80.34 | 0.249 | 0.103 | 0.204 | |
CNNDM | PEFT (reference) | 11.93 | 0.231 | 0.069 | 0.218 |
@@
Avg LoRA | 21.07 | 0.168 | 0.039 | 0.121 | |
TIES | 18.07 | 0.154 | 0.037 | 0.109 | |
DARE | 4.67 | 0.137 | 0.032 | 0.096 | |
Arrow | 13.13 | 0.153 | 0.037 | 0.111 | |
LoraHub | 15.30 | 0.087 | 0.008 | 0.038 | |
MeteoRA (T1-1k) | 11.93 | 0.233 | 0.070 | 0.218 | |
MeteoRA (T2-1k) | 11.93 | 0.232 | 0.070 | 0.219 | |
MeteoRA (T2-100) | 21.11 | 0.205 | 0.054 | 0.146 | |
MeteoRA (T2-5) | 6.52 | 0.203 | 0.054 | 0.143 | |
LingPuzz | PEFT (reference) | 44.12 | 0.785 | 0.589 | 0.734 |
@@
Avg LoRA | 28.87 | 0.421 | 0.134 | 0.331 | |
TIES | 34.17 | 0.432 | 0.134 | 0.339 | |
DARE | 56.23 | 0.357 | 0.113 | 0.281 | |
Arrow | 59.00 | 0.721 | 0.505 | 0.659 | |
LoraHub | 39.28 | 0.245 | 0.063 | 0.184 | |
MeteoRA (T1-1k) | 41.72 | 0.695 | 0.451 | 0.636 | |
MeteoRA (T2-1k) | 41.72 | 0.696 | 0.448 | 0.639 | |
MeteoRA (T2-100) | 50.81 | 0.666 | 0.408 | 0.588 | |
MeteoRA (T2-5) | 46.17 | 0.655 | 0.394 | 0.580 | |
NewsDE | PEFT (reference) | 97.65 | - | - | - |
@@
Avg LoRA | 63.56 | - | - | - | |
TIES | 46.47 | - | - | - | |
DARE | 36.60 | - | - | - | |
Arrow | 37.36 | - | - | - | |
LoraHub | 11.87 | - | - | - | |
MeteoRA (T1-1k) | 86.48 | - | - | - | |
MeteoRA (T2-1k) | 86.48 | - | - | - | |
MeteoRA (T2-100) | 51.42 | - | - | - | |
MeteoRA (T2-5) | 86.48 | - | - | - | |
NewsES | PEFT (reference) | 81.54 | - | - | - |
@@
Avg LoRA | 31.18 | - | - | - | |
TIES | 30.55 | - | - | - | |
DARE | 17.61 | - | - | - | |
Arrow | 31.82 | - | - | - | |
LoraHub | 0.0 | - | - | - | |
MeteoRA (T1-1k) | 81.54 | - | - | - | |
MeteoRA (T2-1k) | 81.54 | - | - | - | |
MeteoRA (T2-100) | 81.54 | - | - | - | |
MeteoRA (T2-5) | 63.72 | - | - | - | |
NewsIT | PEFT (reference) | 54.90 | - | - | - |
@@
Avg LoRA | 38.54 | - | - | - | |
TIES | 37.48 | - | - | - | |
DARE | 52.21 | - | - | - | |
Arrow | 38.02 | - | - | - | |
LoraHub | 0.0 | - | - | - | |
MeteoRA (T1-1k) | 54.90 | - | - | - | |
MeteoRA (T2-1k) | 51.83 | - | - | - | |
MeteoRA (T2-100) | 35.22 | - | - | - | |
MeteoRA (T2-5) | 36.78 | - | - | - | |
StratQA | PEFT (reference) | 10.58 | 0.249 | 0.077 | 0.236 |
@@
Avg LoRA | 38.80 | 0.112 | 0.024 | 0.089 | |
TIES | 10.90 | 0.102 | 0.022 | 0.082 | |
DARE | 14.78 | 0.128 | 0.027 | 0.100 | |
Arrow | 12.19 | 0.206 | 0.057 | 0.165 | |
LoraHub | 14.35 | 0.147 | 0.033 | 0.116 | |
MeteoRA (T1-1k) | 10.58 | 0.252 | 0.076 | 0.239 | |
MeteoRA (T2-1k) | 10.58 | 0.250 | 0.077 | 0.239 | |
MeteoRA (T2-100) | 20.56 | 0.228 | 0.065 | 0.174 | |
MeteoRA (T2-5) | 11.67 | 0.213 | 0.055 | 0.162 | |
TopChat | PEFT (reference) | 39.50 | 0.151 | 0.025 | 0.141 |
@@
LoRA-B | 19.22 | 0.139 | 0.019 | 0.131 | |
Avg LoRA | 23.59 | 0.094 | 0.012 | 0.078 | |
TIES | 26.13 | 0.092 | 0.011 | 0.077 | |
DARE | 38.31 | 0.086 | 0.008 | 0.066 | |
Arrow | 35.64 | 0.112 | 0.016 | 0.091 | |
LoraHub | 0.08 | 0.049 | 0.002 | 0.031 | |
MeteoRA (T1-1k) | 45.64 | 0.152 | 0.026 | 0.141 | |
MeteoRA (T2-1k) | 45.64 | 0.152 | 0.024 | 0.141 | |
MeteoRA (T2-100) | 27.36 | 0.129 | 0.021 | 0.107 | |
MeteoRA (T2-5) | 40.86 | 0.130 | 0.018 | 0.109 |
@@ A.5 Composite-n evaluation results details
LogDeduc | 0.500 | 0.430 | - | - | - | - | - | - | - | - |
QuestSel | 0.545 | 0.630 | - | - | - | - | - | - | - | - |
StratQA | 0.445 | 0.250 | 15.31 | 10.55 | 0.195 | 0.135 | 0.052 | 0.027 | 0.182 | 0.128 |
@@
LogDeduc | 0.500 | 0.500 | - | - | - | - | - | - | - | - |
QuestSel | 0.620 | 0.770 | - | - | - | - | - | - | - | - |
AbsNarr | 0.350 | 0.460 | - | - | - | - | - | - | - | - |
GSWiki | 0.650 | 0.410 | - | - | - | - | - | - | - | - |
StratQA | 0.495 | 0.275 | 9.86 | 9.41 | 0.221 | 0.219 | 0.069 | 0.063 | 0.207 | 0.208 |
@@
LogDeduc | 0.500 | 0.453 | - | - | - | - | - | - | - | - |
QuestSel | 0.703 | 0.688 | - | - | - | - | - | - | - | - |
AbsNarr | 0.625 | 0.672 | - | - | - | - | - | - | - | - |
GSWiki | 0.773 | 0.727 | - | - | - | - | - | - | - | - |
@@
StratQA | 0.461 | 0.211 | 3.23 | 0.00 | 0.225 | 0.106 | 0.051 | 0.025 | 0.210 | 0.099 |
DisflQA | 0.266 | 0.117 | - | - | - | - | - | - | - | - |
NewsDE | - | - | 14.78 | 14.54 | - | - | - | - | - | - |